# A Closed-Form Diffusion Model for Learnring Dynamics from Marginal Observations

### v1

@@ Abstract
Score-based generative models (SGMs) have recently gained significant attention for their strong performance in generating high-fidelity data. These models learn mappings from a simple Gaussian distribution to a complex data distribution. To extend this idea to transformations between arbitrary probability distributions, recent work has considered the Schrödinger Bridge (SB) problem. However, the SB problem rarely admits a closed-form solution and is typically addressed using iterative stochastic simulation methods, which can be unstable and computationally expensive. In this work, we introduce a closed-form framework for learning the stochastic dynamics of SB systems and provide a unified perspective that includes previously known closed-form solutions—such as the Schrödinger–Föllmer process and the Gaussian SB—as special cases. Notably, the classical Gaussian SB solution, previously derived using substantially more sophisticated tools such as Riemannian geometry and generator theory, follows directly from our formulation as an immediate corollary. Building on this foundation, we develop a simulation-free algorithm that infers SB dynamics directly from samples of the source and target distributions. We demonstrate the versatility of our approach in two domains: modeling developmental trajectories in single-cell genomics and performing image restoration tasks including inpainting and deblurring. This framework opens new directions for efficient, scalable nonlinear diffusion modeling across scientific and machine learning applications.
@@ 1 Introduction
Diffusion-based generative models have been extensively developed over the past decade and have found wide-ranging applications in machine learning and related fields, including computer vision (Sohl-Dickstein et al., 2015; Endo and Kanamori, 2020), image analysis (Wang and Ponce, 2021; Pu et al., 2015), natural language processing (Shah, 2021; Fu, 2019; Miao, 2017), drug discovery (Chen et al., 2023; Zeng et al., 2022), and recommender systems (Liu et al., 2021). These models aim to transform a simple, tractable reference distribution—typically a multivariate Gaussian—into a complex data distribution by learning a stochastic process that continuously morphs one into the other. A prominent subclass is the score-based generative models (SGMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Song et al., 2021), which have achieved state-of-the-art performance in generating high-fidelity samples across multiple domains. SGMs are built around two coupled stochastic processes: a forward (diffusion) process and a reverse (denoising) process. The forward process incrementally corrupts clean data by adding Gaussian noise over time, eventually mapping it to a simple prior distribution. The reverse process then learns to invert this corruption, progressively refining pure noise into realistic samples that follow the data distribution. Both directions can be formulated as stochastic differential equations (SDEs), with drift and diffusion terms designed so that the marginal distributions at the endpoints coincide with the data and prior distributions, respectively. The dynamics of the reverse process are governed by the score function—the gradient of the log-probability density—which directs the system toward regions of high data likelihood. In practice, the score function is learned by training a neural network to estimate it at each time step, typically through a denoising score-matching objective. Once trained, the model generates samples by numerically integrating the learned reverse SDE, allowing efficient and stable synthesis of complex data distributions.
@@
While highly successful in data synthesis, SGMs are less directly applicable to many biological and scientific settings—such as single-cell studies—where the goal is not to generate data from noise, but to learn the underlying stochastic dynamics that interpolate between two observed distributions. In these applications, one seeks to infer the most likely probabilistic paths connecting an initial and a terminal state distribution, rather than reconstructing samples from a fixed prior. To address this limitation, a growing body of work has turned to the Schrödinger Bridge (SB) framework (Bernton et al., 2019; Chen et al., 2023; Bortoli et al., 2023; Shi et al., 2023; Chen et al., 2023; Liu et al., 2023; Stromme, 2023), which can be viewed as a natural generalization of diffusion-based models. Originating from Schrödinger’s work in the 1930s (Chetrite et al., 2021; Schrödinger, 1932), SB aims to construct a stochastic process that evolves between two arbitrary marginal distributions while remaining close—in a Kullback–Leibler sense—to a chosen reference process, typically Brownian motion. Equivalently, the SB problem can be formulated as an entropically regularized optimal transport problem, in which the objective is to identify the most likely stochastic transport plan between the two distributions over a finite time horizon.
In practice, however, solving SB problems remains computationally challenging. Except in a few special cases—such as when both endpoint distributions $\mu_{0}$ and $\mu_{1}$ are Gaussian (Bunne et al., 2023) or when $\mu_{0}$ is a Dirac delta—the SB problem generally admits no closed-form solution. Consequently, most approaches rely on iterative numerical approximations that simulate the underlying stochastic process (Bortoli et al., 2023; Chen et al., 2023). While theoretically well-founded, these methods often suffer from numerical instability and poor scalability in high-dimensional spaces (Shi et al., 2023). A widely used approach is the Iterative Proportional Fitting (IPF) algorithm—also known as the Sinkhorn algorithm—which alternates between solving forward and backward “half-bridges”. At each iteration, the drift term of the associated SDE is re-estimated via regression on samples generated from the previous half-bridge, typically requiring the training of deep neural networks. As a result, the drift at each iteration depends on the cumulative sum of scores estimated across all earlier iterations, and each step entails storing multiple neural networks (usually two per iteration). This iterative training and storage burden makes SB methods computationally expensive and difficult to scale, especially for high-dimensional generative modeling tasks.
@@
To overcome these computational limitations, we propose a closed-form framework for learning the stochastic dynamics between two arbitrary distributions. We derive, for the first time, an explicit analytical relation between the dynamic formulation of the Schrödinger Bridge (SB) problem and the solution to its corresponding static problem. Specifically, we show that the solutions of the SB system admit analytic expressions that depend solely on integrations with respect to the endpoint joint distributions. Remarkably, this result yields a unifying perspective that encompasses all previously known closed-form SB solutions, including the Schrödinger–Föllmer process—which transports a degenerate Dirac delta distribution at time zero to a target distribution at time one—and the Gaussian Schrödinger Bridge, which transports one multivariate Gaussian to another, as special cases.
@@
The remainder of the paper is organized as follows. Section 2 introduces the Schrödinger Bridge diffusion process and presents a tractable formulation that can be used to derive explicit closed-form solutions to the Schrödinger systems. Section 3 develops two algorithms: a training-free algorithm (Algorithm 1) in Section 3.1, derived directly from our main theorem, and a simulation-free algorithm (Algorithm 2) in Section 3.2, based on an equivalent variational representation of the same result. Section 4 presents numerical experiments evaluating the performance of our methods on both simulated low-dimensional data and real-world high-dimensional image datasets. Finally, Section 5 provides concluding remarks, and Appendix contains proofs of the main theoretical results.
@@ Appendix D Proof of Theorem 2
$\mathbb{E}_{{\bf Q}_{{\bf x}_{0},{\bf x}_{1},{\bf x}_{t}}}\left[\left|\sigma(t)^{2}{\bf s}_{{\bf x}_{1}}(t,{\bf x}_{t})-{\bf u}({\bf x}_{t},t)\right|^{2}\right]$
@@
$\geq$ $\mathbb{E}_{{\bf Q}_{{\bf x}_{0},{\bf x}_{1},{\bf x}_{t}}}\left[\left|\sigma(t)^{2}{\bf s}_{{\bf x}_{1}}(t,{\bf x}_{t})-{\bf u}^{\star}({\bf x}_{t},t)\right|^{2}\right].$ (D.1)

### v2

# A Closed-Form Framework for Schrödinger Bridges Between Arbitrary Densities

@@ Abstract
Score-based generative models have recently attracted significant attention for their ability to generate high-fidelity data by learning maps from simple Gaussian priors to complex data distributions. A natural generalization of this idea to transformations between arbitrary probability distributions leads to the Schrödinger Bridge (SB) problem. However, SB solutions rarely admit closed-form expressions and are commonly obtained through iterative stochastic simulation procedures, which are computationally intensive and can be unstable. In this work, we introduce a unified closed-form framework for representing the stochastic dynamics of SB systems. Our formulation subsumes previously known analytical solutions—including the Schrödinger–Föllmer process and the Gaussian SB—as specific instances. Notably, the classical Gaussian SB solution, previously derived using substantially more sophisticated tools such as Riemannian geometry and generator theory, follows directly from our formulation as an immediate corollary. Leveraging this framework, we develop a simulation-free algorithm that infers SB dynamics directly from samples of the source and target distributions. We demonstrate the versatility of our approach in two settings: (i) modeling developmental trajectories in single-cell genomics and (ii) solving image restoration tasks such as inpainting and deblurring. This work opens a new direction for efficient and scalable nonlinear diffusion modeling across scientific and machine learning applications.
@@ 1 Introduction
Diffusion-based generative models have seen rapid development and broad application across machine learning domains, including computer vision (Sohl-Dickstein et al., 2015), image analysis (Wang and Ponce, 2021; Pu et al., 2015), natural language processing (Shah, 2021; Fu, 2019), and drug discovery (Chen et al., 2023). These models construct a stochastic process that transforms a simple reference distribution, such as a multivariate Gaussian, into a complex data distribution. A central class within this family is score-based generative models (SGMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Song et al., 2021), which currently achieve state-of-the-art sample quality across multiple domains. SGMs couple two stochastic processes: a forward diffusion process that gradually corrupts data into noise, and a reverse denoising process that reconstructs data from noise. Both can be described as stochastic differential equations whose endpoints match the data and prior distributions.
@@
While SGMs excel at data synthesis, many scientific applications—such as single-cell analysis—require learning the stochastic dynamics that interpolate between two observed distributions, rather than generating samples from noise. In these settings, the goal is to infer the most likely probabilistic paths connecting initial and terminal state distributions. This has led to growing interest in the Schrödinger Bridge (SB) framework (Bernton et al., 2019; Chen et al., 2023; Bortoli et al., 2023; Shi et al., 2023; Chen et al., 2023; Liu et al., 2023; Stromme, 2023), which generalizes diffusion-based models. SB seeks a stochastic process that evolves between two marginals while staying close to a reference process, typically Brownian motion.
@@
In general, SB problems lack closed-form solutions except in special cases (e.g., both marginals Gaussian (Bunne et al., 2023)). As a result, most methods rely on iterative numerical approximations (Bortoli et al., 2023; Chen et al., 2023), which can be unstable and scale poorly in high dimensions (Shi et al., 2023). To overcome these computational limitations, we propose a closed-form framework for learning the stochastic dynamics between two arbitrary distributions. We derive, for the first time, an explicit analytical relation between the dynamic formulation of the SB problem and the solution to its corresponding static problem. Specifically, we show that the solutions of the SB system admit analytic expressions that depend solely on integrations with respect to the endpoint joint distributions. Remarkably, this result yields a unifying perspective that encompasses all previously known closed-form SB solutions, including the Schrödinger–Föllmer process—which transports a degenerate Dirac delta distribution at time zero to a target distribution at time one—and the Gaussian Schrödinger Bridge, which transports one multivariate Gaussian to another, as special cases.
@@
The remainder of the paper is organized as follows. Section 2 introduces the Schrödinger Bridge diffusion process and presents a tractable formulation that can be used to derive explicit closed-form solutions to the Schrödinger systems. Section 3 develops two algorithms: a training-free algorithm (Algorithm 1) in Section 3.1, derived directly from our main theorem, and a simulation-free algorithm (Algorithm 2) in Section 3.2, based on an equivalent variational representation of the same result. Section 4 presents numerical experiments evaluating the performance of our methods on both simulated low-dimensional data and real-world high-dimensional image datasets. Finally, Section 5 provides concluding remarks, and the proofs of the main theoretical results are included in the supplimentary material.
@@ Appendix D Proof of Theorem 2
$\mathbb{E}_{{\bf Q}_{{\bf x}_{0},{\bf x}_{1},{\bf x}_{t}}}\left[\left|\sigma(t)^{2}{\bf s}_{{\bf x}_{1}}(t,{\bf x}_{t})-{\bf u}({\bf x}_{t},t)\right|^{2}\right]$ (D.1)
@@
$\geq$ $\mathbb{E}_{{\bf Q}_{{\bf x}_{0},{\bf x}_{1},{\bf x}_{t}}}\left[\left|\sigma(t)^{2}{\bf s}_{{\bf x}_{1}}(t,{\bf x}_{t})-{\bf u}^{\star}({\bf x}_{t},t)\right|^{2}\right].$